OpenStudy (anonymous):

I need help with: SquareRoot 12 + 3 SquareRoot 75 = and SquareRoot 20 + 2 SquareRoot 45 - 3 SquareRoot 80 = Or any guidance of how to do them?. Thanks

OpenStudy (anonymous):

\[\sqrt{12} + 3\sqrt{75}\] and \[\sqrt{20} + 2\sqrt{45} - 3\sqrt{80}\]

OpenStudy (anonymous):

\[12=4\times 3\] so \[\sqrt{12}=\sqrt{4\times 3}=\sqrt{4}\times \sqrt{3}=2\sqrt{3}\]

OpenStudy (anonymous):

similarly \[75=25\times 3\] and so \[\sqrt{75}=5\sqrt{3}\]

OpenStudy (anonymous):

s o your first question is \[sqrt{12}+3\sqrt{75}=2\sqrt{3}+3\times 5\sqrt{3}=2\sqrt{3}+15\sqrt{3}=18\sqrt{3}\]

OpenStudy (anonymous):

its pretty easy, just check if perfect squares are factors of the think under the sqrt

OpenStudy (anonymous):

actually 2 + 15 = 17 doesn't it?

OpenStudy (anonymous):

1,4,9,16,25,36 ... etc

OpenStudy (anonymous):

like for sqrt (80) 16 x 5 = 80

OpenStudy (anonymous):

so i suppose unless i made some other mistake \[17\sqrt{3}\] is it

OpenStudy (anonymous):

and 16 is 4^2

OpenStudy (anonymous):

so sqrt(80) = 4 sqrt(5)

OpenStudy (anonymous):

Thank You very much for the explanation satellite and also thank you elecengineer !

OpenStudy (anonymous):

for the second question sqrt(20) 4 is a factor of 20 so that = sqrt(2^2 x 5 ) = sqrt(2^2) x sqrt(5) = 2sqrt(5)

OpenStudy (anonymous):

similairly with 45, 9 is a factor , which is 3^2